# Compressed data structures for Heegaard splittings

### v1

@@ Appendix A Extension to general systems
As mentioned in Remark 2.1, there is a generalization that allows systems to have more than $g$ components. In particular, for this section, we will assume that a system is a multicurve of cardinality at least $g$ and with no isotopic components. A standard Euler characteristic argument [50] implies that a system $\gamma$ cuts $\Sigma_{g}$ into #$\#\gamma-g+1$ at least thrice punctured spheres. When #$\#\gamma=g$, we say that the system is minimal and it reduces to the definition used in the main text. If, on the other hand, #$\#\gamma=3g-3$, we say that the system is maximal or a pants decomposition because it cuts $\Sigma_{g}$ into a union of three-times punctured spheres known as pairs of pants.
Like minimal systems, general systems uniquely define handlebodies. This means that the notion of equivalence naturally extends to the more general systems, even if they have different cardinality. In particular, one [16, 31] can show that, provided that $\gamma$ is not a pants decomposition, the band sum $c$ of two punctures $c_{1}$ and $c_{2}$ and an arc $p$ connecting them in $\Sigma_{g}\backslash\gamma$ is a meridian of $\mathcal{H}_{\gamma}$ and is not isotopic to any other component of $\gamma$. This means that $\gamma^{\prime}=\gamma\cup\{c\}$ forms a system which is equivalent to $\gamma$; we call $\gamma^{\prime}$ an extension of $\gamma$.
We can expand our data structure to allow for more general systems for $\alpha$ and $\beta$. In particular, a diagram $(\Sigma_{g},\alpha,\beta)$ is minimal if $\alpha$ and $\beta$ are minimal systems; similarly it is maximal if they are pants decomposition. In fact, although more resource-intensive, this general version of Heegaard diagrams has been used both in mathematical (for example, [25, 56]) and in computer science ([16]) settings. All algorithms described in this paper can be adapted to this more general notion of Heegaard diagrams with only minor modifications. We can, moreover, describe efficient methods to translate between minimal and more general systems. This is exactly the content of the next two results.
@@
Compute $S(T,\beta)$ and delete the separating components of $\beta$ by using breadth-first in the dual graph. In total, this takes time $O(m|T|+g|T|)=O(m|T|)$ since $O(g)=O(|T|)=O(\sqrt{m})$. Similarly, delete separating curves in $\alpha$ in time $O(g|T|)$. ∎
@@
Note that $\Sigma_{g}\backslash\gamma$ is homemorphic to the $2g$ punctured sphere, so let $d_{1},\dots,d_{2g}$ represent the boundary components of this surface. We will find some $32g-$ 3 new separating curves $d_{2g+1},\dots,d_{3g-3}$ in $\Sigma_{g}$ that recursively define pairs of pants as follows. In step $i$, choose a curve $d_{g+j},0\leq j\leq i-1$ from a list of available curves. Call this choice $s_{1}$ and pick a face $t$ adjacent to $s_{1}$. Using breadth-first search, we find, in time $O(|T|)$, the shortest path $p$ in the dual graph starting from $t$ to any face adjacent to one of the available curves. Call the closest available curve $s_{2}$ and remove $s_{1}$ and $s_{2}$ from the list of available curves. Note that we can encode $p$ as a list of faces of $T$. In time $O(\|E_{T}(s_{1})\|)=O(|T|)$, we update $p$ by iteratively removing all faces adjacent to $s_{1}$ except for the last one. By construction, $p$ intersects each edge of the cellular complex at most once and does not intersect any face adjacent to an available curve, except for exactly one face adjacent to $E_{T}({s_{1}})$ and one face adjacent to $E_{T}({s_{2}})$.
As in the proof of Theorem 4.5, make $I_{T}(d_{i})=I_{T}(s_{2})^{-1}\circ I_{T}(p)^{-1}\circ I_{T}(s_{1})\circ I_{T}(p)$, where $p$ is potentially trivial, and add $d_{i}$ to the list of available curves. The curve $d_{i}$ separates a thrice puncture sphere component with punctures defined by $s_{1}$, $s_{2}$, and $d_{i}$, from a $2g-i$ punctured sphere component, and which component is which can be identified in time $O(|T|)$. We then update $T$ to a cellular embedding of $\Sigma_{0}^{2g+i}$ by adding the edges and vertices defined by $I_{T}(d_{i})$ to the graph and blocking paths in the dual graph transverse to them. While $T$ becomes more complicated at each step, because the arc $p$ is constructed to intersect a constant number of faces adjacent to the available curves and the curves $d_{g+j}$, $11\leq j\leq i-$ 1, are separating, at each step, breadth-first needs only to consider $O(|T|)$ many edges and vertices of the dual graph of $T$. The algorithm terminates when $i=3g-3$, in which case we have defined, by construction, a maximal system $\gamma=\{c_{1},\dots,c_{g},d_{g+1},\dots,d_{3g-3}\}$. ∎
@@
Theorem A.2 does not guarantee that there exists a $\phi\in\operatorname{Mod}(\Sigma_{g})$ such that $\beta^{\prime}=\phi(\alpha^{\prime})$ where $\alpha^{\prime}$ and $\beta^{\prime}$ are pants decompositions. In fact, it is a well-known result from surface theory [44] that such a homeomorphism exists if and only if the pants decomposition graphs associated with $\alpha^{\prime}$ and $\beta^{\prime}$ are isotopic. One can easily imagine a modification to the algorithm of the theorem that constructs a pants decomposition $\beta^{\prime}$ that defines a graph isotopic to the graph defined by $\alpha^{\prime}$ by forcing arcs between particular punctures of $\Sigma_{g}\backslash\alpha^{\prime}$ as in Theorem 4.5, but that would imply an algorithm exponential on $|T|$. Alternatively, a method based on the Hatcher and Thurston proof of connectivity of the pants graph (not to be confused with the pants decomposition graph) [23][Appendix] can be devised, but efficiency in $g$ is not trivial.
@@ Appendix B Details on the implementations of Section 5
As mentioned in the main text, several simplifying assumptions were made for the implementation of the examples in Section 5. First and most importantly, we did not compute a data structure for the Heegaard diagrams as described in Section 3. Instead, we use general cellular embedding $T_{g}$ given by the fundamental polygons of $\Sigma_{g}$ [22] to represent the surfaces, which implies that, instead of normal coordinates, we assumed the $\beta$ curves to be given by SLPs $I_{T}(\beta)$. Although this deviates from what was described in this paper, we note that cyclic reducing the SLPs is irrelevant to finding $H_{1}(M)$ and introduces some unbearable computational cost. We also assumed the Lickorish generators to be edged with respect to $T_{g}$ and that their intersection sequences are computed only once, say as a preprocessing step, using Lemma 3.5 (in fact, for these experiments, we computed the SLPs by hand). Figure 8 shows $T_{g}$, for $g=1$ and $2$, while Table 2 indicates the edged and (uncompressed) intersection sequences of the curves involved.
@@
Table 2. Here, we use []delimited-[][$\cdot$][ ⋅ ] to denote lists. We assume that the $\alpha$ curves are contained in the Lickorish generators, meaning that they are omitted in the fourth and fifth columns.

### v2

@@ Appendix A Extension to general systems
As mentioned in Remark 2.1, there is a generalization that allows systems to have more than $g$ components. In particular, for this section, we will assume that a system is a multicurve of cardinality at least $g$ and with no isotopic components. A standard Euler characteristic argument [schleimer2006notes] implies that a system $\gamma$ cuts $\Sigma_{g}$ into $\gamma-g+1$ at least thrice punctured spheres. When $\gamma=g$, we say that the system is minimal and it reduces to the definition used in the main text. If, on the other hand, $\gamma=3g-3$, we say that the system is maximal or a pants decomposition because it cuts $\Sigma_{g}$ into a union of three-times punctured spheres known as pairs of pants.
Like minimal systems, general systems uniquely define handlebodies. This means that the notion of equivalence naturally extends to the more general systems, even if they have different cardinality. In particular, one [ennes2025hardnesscomputationquantuminvariants, johnson2006notes] can show that, provided that $\gamma$ is not a pants decomposition, the band sum $c$ of two punctures $c_{1}$ and $c_{2}$ and an arc $p$ connecting them in $\Sigma_{g}\backslash\gamma$ is a meridian of $\mathcal{H}_{\gamma}$ and is not isotopic to any other component of $\gamma$. This means that $\gamma^{\prime}=\gamma\cup\{c\}$ forms a system which is equivalent to $\gamma$; we call $\gamma^{\prime}$ an extension of $\gamma$.
We can expand our data structure to allow for more general systems for $\alpha$ and $\beta$. In particular, a diagram $(\Sigma_{g},\alpha,\beta)$ is minimal if $\alpha$ and $\beta$ are minimal systems; similarly, it is maximal if they are pants decompositions. In fact, although more resource-intensive, this general version of Heegaard diagrams has been used both in mathematical (for example, [hempel20013, yoshizawa2014high]) and in computer science ([ennes2025hardnesscomputationquantuminvariants]) settings. All algorithms described in this paper can be adapted to this more general notion of Heegaard diagrams with only minor modifications. We can, moreover, describe efficient methods to translate between minimal and more general systems. This is exactly the content of the next two results.
@@
Compute $S(T,\beta)$ and delete the separating components of $\beta$ by using breadth-first search in the dual graph. In total, this takes time $O(m|T|+g|T|)=O(m|T|)$ since $O(g)=O(|T|)=O(\sqrt{m})$. Similarly, delete any separating curve of $\alpha$ in time $O(g|T|)$. ∎
@@
Note that $\Sigma_{g}\backslash\gamma$ is homeomorphic to the $2g$ punctured sphere, so let $d_{1},\dots,d_{2g}$ represent the boundary components of this surface. We will find some 2g32g-3 new separating curves $d_{2g+1},\dots,d_{3g-3}$ in $\Sigma_{g}$ that recursively define pairs of pants as follows. In step $i$, choose a curve $d_{g+j},0\leq j\leq i-1$ from a list of available curves. Call this choice $s_{1}$ and pick a face $t$ adjacent to $s_{1}$. Using breadth-first search, we find, in time $O(|T|)$, the shortest path $p$ in the dual graph starting from $t$ to any face adjacent to one of the available curves. Call the closest available curve $s_{2}$ and remove $s_{1}$ and $s_{2}$ from the list of available curves. Note that we can encode $p$ as a list of faces of $T$. In time $O(\|E_{T}(s_{1})\|)=O(|T|)$, we update $p$ by iteratively removing all faces adjacent to $s_{1}$ except for the last one. By construction, $p$ intersects each edge of the cellular complex at most once and does not intersect any face adjacent to an available curve, except for exactly one face adjacent to $E_{T}({s_{1}})$ and one face adjacent to $E_{T}({s_{2}})$.
As in the proof of Theorem 4.5, make $I_{T}(d_{i})=I_{T}(s_{2})^{-1}\circ I_{T}(p)^{-1}\circ I_{T}(s_{1})\circ I_{T}(p)$, where $p$ is potentially trivial, and add $d_{i}$ to the list of available curves. The curve $d_{i}$ separates a thrice-punctured sphere component with punctures defined by $s_{1}$, $s_{2}$, and $d_{i}$, from a 2gi2g-i punctured sphere component, and which component is which can be identified in time $O(|T|)$. We then update $T$ to a cellular embedding of $\Sigma_{0}^{2g+i}$ by adding the edges and vertices defined by $I_{T}(d_{i})$ to the graph and blocking paths in the dual graph transverse to them. While $T$ becomes more complicated at each step, because the arc $p$ is constructed to intersect a constant number of faces adjacent to the available curves and the curves $d_{g+j}$, $1\leq j\leq i-1$, are separating, at each step, breadth-first needs only to consider $O(|T|)$ many edges and vertices of the dual graph of $T$. The algorithm terminates when i=3g3i=3g-3, in which case we have defined, by construction, a maximal system $\gamma=\{c_{1},\dots,c_{g},d_{g+1},\dots,d_{3g-3}\}$. ∎
@@
Theorem A.2 does not guarantee that there exists a $\phi\in\operatorname{Mod}(\Sigma_{g})$ such that $\beta^{\prime}=\phi(\alpha^{\prime})$ where $\alpha^{\prime}$ and $\beta^{\prime}$ are pants decompositions. In fact, it is a well-known result from surface theory [moser2014upper] that such a homeomorphism exists if and only if the pants decomposition graphs associated with $\alpha^{\prime}$ and $\beta^{\prime}$ are isotopic. One can easily imagine a modification to the algorithm of the theorem that constructs a pants decomposition $\beta^{\prime}$ that defines a graph isotopic to the graph defined by $\alpha^{\prime}$ by forcing arcs between particular punctures of $\Sigma_{g}\backslash\alpha^{\prime}$ as in Theorem 4.5, but that would imply an algorithm exponential on $|T|$. Alternatively, a method based on Hatcher and Thurston’s proof of connectivity of the pants graph (not to be confused with the pants decomposition graph) [hatcher1980presentation][Appendix] can be devised, but efficiency in $g$ is not trivial.
@@ Appendix B Details on the implementations of Section 5
As mentioned in the main text, several simplifying assumptions were made for the implementation of the examples in Section 5 and Appendix 5.2. First and most importantly, we did not compute a data structure for the Heegaard diagrams as described in Section 3. Instead, we use a general cellular embedding $T_{g}$ given by the fundamental polygons of $\Sigma_{g}$ [hatcher2002algebraic] to represent the surfaces, which implies that, instead of normal coordinates, we assume the $\beta$ curves to be given by SLPs $I_{T}(\beta)$. Although this deviates from what was described in this paper, we note that cyclically reducing the SLPs is irrelevant to finding $H_{1}(M)$ (or $\pi_{1}(M)$) and introduces some unbearable computational cost. We also assumed the Lickorish generators to be edged with respect to $T_{g}$ and that their intersection sequences are computed only once, say as a preprocessing step, using Lemma 3.4 (in fact, for these experiments, we computed the SLPs by hand). Figure 12 shows $T_{g}$, for $g=1$ and $2$, while Table 2 indicates the edges and (uncompressed) intersection sequences of the curves involved.
@@
Table 2. Here, we use $\cdot$] to denote lists. We assume that the $\alpha$ curves are contained in the Lickorish generators, meaning that they are omitted in the fourth and fifth columns.